the sum of the digits of a two digit number is 12, if the digits are reversed, the new number is 18 more than the original number. what is the original number?
Answer by mananth(16946)

let x be in the tens place
y in the units place
the number will be (10x+y)
x+y =12 .....(1) The sum of the digits
10y+x=10x+y +18
9y-9x=18
/9
y-x=2... (2)
add (1) and (2)
2y=14
y=7
plug y in (1)
x=5
the number is 57